# Exact State Evolution and Energy Spectrum in Solvable Bosonic Models

### v1

@@ Abstract
Solvable bosonic models provide a fundamental framework for describing light propagation in nonlinear media, including optical down-conversion processes that generate squeezed states of light and their higher-order generalizations. In quantum optics a central objective is to determine the time evolution of a given initial state. Exact analytic solution to the state-evolution problem is presented, applicable to a broad class of solvable bosonic models and arbitrary initial states. Moreover, the characteristic equation governing the energy spectrum is derived and the eigenstates are found in the form of continued fractions and as principal minors of the associated Jacobi matrix. The results provide a solid analytical framework for discussion of exactly solvable bosonic models.
@@ I Introduction
The structure of the paper is as follows. In Section II, the class of bosonic models to which the present approach applies is outlined. Section III provides the solution of the state-evolution problem for an arbitrary initial state by evaluating the average of the unitary evolution operator between two arbitrary basis states within each invariant subspace of the Hilbert space. Theorem1 and Corollaries1 and 2 summarize the main results of this section. In Section IV, the characteristic polynomial determining the energy eigenvalues in each invariant subspace is derived, and the quantum amplitudes of the corresponding eigenstates are expressed in terms of continued fractions and as Jacobi determinants. Theorem 2 and Corollary 3 contain the main results concerning the energy spectrum. Furthermore, the stationary state (the eigenstate with zero energy) is obtained in explicit analytic form. Section V discusses the importance of the results and their possible use in applications. Finally, Section VI presents a brief summary of the results and points on an open problem.
@@ III.3 The $g$-factors as powers of a Hessenberg matrix
The matrix $\mathbf{B}$ of rank $N$ belongs to the class of lower Hessenberg matrices (only the main diagonal and the first superdiagonal may be nonzero). Powers of such matrices are easy to compute numerically Hessen . The $l$th power of $\mathbf{B}$ admits the representation
@@
(the elements in the lower tringular part are all equal to $1$), we get an equivalent expression
@@ III.4 The state evolution: main results
where the orthonormal vectors $|e_{m}\rangle$, $m=0,\ldots,N$ span an auxiliary vector space (isomorphic to the considered subspace $\mathcal{H}_{N}$ of the Hilbert space), the lower-Hessenberg matrix $\mathbf{B}$ is defined in Eq. (35) and the lower-triangular matrix $\mathbf{T}$ in Eq. (38). Here $l_{0}(k,n)=\mathrm{max}(0,k-n)$, i.e., for $n<k$ the nonzero terms have $l\geq k-n$. Alternatively, by using Eq. (43), Eq. (2) can be rewritten as
@@ IV The energy spectrum
One approach to determining the energy spectrum of a solvable model is to use the algebraic or group-theoretic methods GroupMeth ; LieAlg3Boson ; GroupMeth2 ; LieAlgGen1 ; LieAlgGen2 . Here an alternative approach is developed. The characteristic polynomial for the energy eigenvalues is derived from a recursion, where the coefficients are expressed as nested sums. Alternatively, the characteristic polynomial can be represented by a continued fraction. The amplitudes of the eigenstates are given in several equivalent forms: as continued fractions (or, equivalently, as a product of the Möbius group matrices in two-dimensional projective space), as nested sums, and as principal minors of a Jacobi matrix.
Consider the eigenvalue problem for the Hamiltonian of Eq. (2) in an invariant subspace $\mathcal{H}_{N}$. Expanding eigenstates in the basis of Eq. (4),
@@
IV.1 The spectrum via continuous fractions
@@
The characteristic equation for the eigenvalues of Eq. (49) reads: $\psi_{N+1}(\lambda)=0$, where $\psi_{n}(\lambda)$ is the general solution. By Eq. (65) this implies $Y_{N+1}(\lambda)=0$, since the denominator on the right-hand side of Eq. (65) is zero for $n=N+1$ (recall that $\beta_{N}=0$).
@@ IV.2 The characteristic equation for the energy eigenvalues
The recursion in Eq. (63), formally considered for unbounded $n\geq 0$, defines a sequence of polynomials in $\lambda$: $Y_{n}=Y_{n}(\lambda)$. It can be easily solved in a similar was as the recursion in Eq. (28). We have the following result.
@@
Now let us prove that the polynomial of order $n+1$ on the right hand side of Eq. (66) solves the recursion in Eq. (63). First of all, Eq. (66) reproduces the first two polynomials: $Y_{0}=1$ and $Y_{1}=\lambda$ (for $n=0$ and $n=1$ there is only the term with $l=0$). Now consider the expression for $Y_{n+1}(\lambda)$ provided by Eq. (66). By separating the term with $l=0$ and using the identity (67), we obtain:
@@ IV.4 The stationary state
The products in Eq. (IV.4) can be recast in this case as follows:
@@
In this case from Eq. (IV.4) one obtains the following (un-normalized) stationary state in the invariant subspace $\mathcal{H}_{2M}$
@@ V Discussion
On the other hand, perturbative series expansions involve powers of boson creation and annihilation operators, i.e., operators with unbounded norms, when one adopts the so-called parametric approximation for the strong pump mode. Such perturbative series lack well-defined domains of convergence in the rigorous mathematical sense, which limits their reliability—one is left to verify their validity only by comparison with experimental results. A prominent example of this is the spontaneous down-conversion process in quadratic nonlinear media and its generalizations to higher-order nonlinearities GenSqueez . The standard perturbative approach based on the parametric approximation has been shown to be deficient for higher-order processes, as it leads to a divergence of the system state at a finite interaction time if the pump field is treated not as a separate quantum degree of freedom but merely as a parameter Buzek93 ; Drobny94 . It is, in fact, by chance that no divergence occurs in the quadratic case, where the standard parametric approximation yields the squeezed state of light ReviewSq , which remains regular in norm. Nevertheless, the domain of validity of the standard parametric approximation had long remained undefined. Only very recently has this issue been addressed through direct comparison with a non-perturbative, quasi-exact approach ourPRA . The latter was constructed upon the existence of an exact solution previously derived for the evolution of a special initial state in Ref. My1 .
@@
The limitations of the current approach stem from its very generality: since the solution is fully characterized by a single polynomial parameter, $\beta_{n}$ in Eq.(4) of Section II, distinct physical regimes correspond to different choices of this polynomial (see Ref. My1 for details), even though the form of the solution remains the same. Thus, to advance the understanding of specific models, one must go beyond generality and focus on particular choices of $\beta_{n}$ representing distinct physical situations. The significance of the present work lies in providing a rigorous, divergence-free framework for such future analyses of particular models.
@@ VI Conclusion
The explicit solution to the arbitrary state evolution problem has been found in the form of an infinite time-series expansion whose coefficients are nested sums involving this polynomial. The characteristic polynomial for the energy eigenvalues has coefficients of a similar structure. The eigenstate amplitudes have been represented equivalently as continued fractions, as products of Möbius-group matrices in the projective space, and as principal minors of the associated Jacobi matrix. The results have important area of applications to the nonlinear quantum-optical models, including $k$-photon down-conversion, which play a central role in modern quantum technologies.
@@
where $P_{\mu}$ is the permutation operator acting on the product of $m$ factors, where $\mu$ runs over different orderings of $m$ factors, with $l$ factors of one type and mlm-l of another type (there are $\binom{m}{l}$ of different orderings). Using Eq. (13) and (14) we get from Eq. (81):

### v2

@@ Abstract
Solvable bosonic models provide a fundamental framework for describing light propagation in nonlinear media, including optical down-conversion processes that generate squeezed states of light and their higher-order generalizations. In quantum optics a central objective is to determine the time evolution of a given initial state. Exact analytic solution to the state-evolution problem is presented, applicable to a broad class of solvable bosonic models and arbitrary initial states. Moreover, the characteristic equation governing the energy spectrum is derived and the eigenstates are found in the form of continued fractions and as the principal minors of the associated Jacobi matrix. The results provide a solid analytical framework for discussion of exactly solvable bosonic models.
@@ I Introduction
The structure of the paper is as follows. In Section II, the class of bosonic models to which the present approach applies is outlined. Section III provides the solution of the state-evolution problem for an arbitrary initial state by giving the average of the unitary evolution operator between two arbitrary basis states within each invariant subspace of the Hilbert space. Theorem1 and Corollaries1 and 2 summarize the main results of this section. In Section IV, the characteristic polynomial determining the energy eigenvalues in each invariant subspace is derived, and the quantum amplitudes of the corresponding eigenstates are expressed in terms of continued fractions and as Jacobi determinants. Theorem 2 and Corollary 3 contain the main results concerning the energy spectrum. Furthermore, the stationary state (the eigenstate with zero energy) is obtained in explicit analytic form. Section V discusses the importance of the results and their possible use in applications. Finally, Section VI presents a brief summary of the results and points on an open problem.
@@ III.3 The $g$-factors as powers of a Hessenberg matrix
The matrix $\mathbf{B}$ of rank $N$ belongs to the class of lower Hessenberg matrices (in contrast to the lower triangular matrix, the first super-diagonal can be nonzero also). Powers of such matrices are easy to compute numerically Hessen . The $l$th power of $\mathbf{B}$ admits the representation
@@
(the elements in the lower triangular part are all equal to $1$), we get an equivalent expression
@@ III.4 The state evolution: main results
where the orthonormal vectors $|e_{m}\rangle$, $m=0,\ldots,N$ span an auxiliary vector space (isomorphic to the considered subspace $\mathcal{H}_{N}$ of the Hilbert space), the lower Hessenberg matrix $\mathbf{B}$ is defined in Eq. (35) and the lower-triangular matrix $\mathbf{T}$ in Eq. (38). Here $l_{0}(k,n)=\mathrm{max}(0,k-n)$, i.e., for $n<k$ the nonzero terms have $l\geq k-n$. Alternatively, by using Eq. (43), Eq. (2) can be rewritten as
@@ IV The energy spectrum
One approach to determining the energy spectrum of a solvable model is to use the algebraic or group-theoretic methods GroupMeth ; LieAlg3Boson ; GroupMeth2 ; LieAlgGen1 ; LieAlgGen2 . Here an alternative approach is developed. The characteristic polynomial for the energy eigenvalues is derived from a recursion, where the coefficients are expressed as nested sums. Alternatively, the characteristic polynomial can be represented by a continued fraction. The amplitudes of the eigenstates are given in several equivalent forms: as continued fractions (or, equivalently, as a product of the Möbius group matrices in two-dimensional projective space), as nested sums, and as the principal minors of a Jacobi matrix.
Consider the eigenvalue problem in an invariant subspace $\mathcal{H}_{N}$ for the Hamiltonian of Eq. (2). Expanding eigenstates in the basis of Eq. (4),
@@
IV.1 The spectrum via continued fractions
@@
The characteristic equation for the eigenvalues of Eq. (49) reads: $\psi_{N+1}(\lambda)=0$, where $\psi_{n}(\lambda)$ is the general solution. We have shown that it also reads $Y_{N+1}(\lambda)=0$ Eq. (64). Observe that the relation between the two forms by Eq. (65) for $n=N+1$ involves the singular first factor on the right-hand side (recall that $\beta_{N}=0$).
@@ IV.2 The characteristic equation for the energy eigenvalues
The recursion in Eq. (63) defines a sequence of polynomials in $\lambda$: $Y_{n}=Y_{n}(\lambda)$. It can be easily solved in a similar way as the recursion in Eq. (28). We have the following result.
@@
Now let us prove that the polynomial of degree $n+1$ on the right hand side of Eq. (66) solves the recursion in Eq. (63). First of all, Eq. (66) reproduces the first two polynomials: $Y_{0}=1$ and $Y_{1}=\lambda$ (for $n=0$ and $n=1$ there is only the term with $l=0$). Now consider the expression for $Y_{n+1}(\lambda)$ provided by Eq. (66). By separating the term with $l=0$ and using the identity (67), we obtain:
@@ IV.4 The stationary state
The products in Eq. (75) can be recast in this case as follows:
@@
In this case from Eq. (75) one obtains the following (un-normalized) stationary state in the invariant subspace $\mathcal{H}_{2M}$
@@ V Discussion
On the other hand, perturbative series expansions involve powers of boson creation and annihilation operators, i.e., operators with unbounded norms, when one adopts the so-called parametric approximation for the strong pump mode. Such perturbative series lack well-defined domains of convergence in the rigorous mathematical sense, which limits their reliability—one is left to verify their validity only by comparison with experimental results. A prominent example of this is the spontaneous down-conversion process in quadratic nonlinear media and its generalizations to higher-order nonlinearities GenSqueez . The standard perturbative approach based on the parametric approximation has been shown to be deficient for higher-order processes, as it leads to a divergence of the system state at a finite interaction time if the pump field is treated not as a separate quantum degree of freedom but merely as a parameter Buzek93 ; Drobny94 . It is, in fact, by chance that no divergence occurs in the quadratic case, where the standard parametric approximation yields the squeezed state of light ReviewSq , which remains regular in norm. Nevertheless, the domain of validity of the standard parametric approximation had long remained undefined. Only very recently has this issue been addressed through direct comparison with a non-perturbative asymptotic approach ourPRA . The latter was constructed upon the existence of an exact solution previously derived for the evolution of a special initial state in Ref. My1 .
@@
The limitations of the current approach stem from its very generality: since the solution is fully characterized by a single polynomial parameter, $\beta_{n}$ in Eq.(4) of Section II, distinct physical regimes correspond to different choices of this polynomial (see Ref. My1 for details), even though the form of the solution remains the same. Thus, to advance the understanding of specific models, one must go beyond generality and focus on a particular $\beta_{n}$ representing a distinct physical situation. Clearly, this task requires development of additional approaches, supplementing the present general framework. The significance of the present work lies in providing a rigorous, divergence-free framework for such future analyses of particular models.
@@ VI Conclusion
The explicit solution to the arbitrary state evolution problem has been found in the form of an infinite time-series expansion whose coefficients are nested sums involving the characterizing polynomial. The characteristic polynomial for the energy eigenvalues has coefficients of a similar structure. The eigenstate amplitudes have been represented equivalently as continued fractions, as products of Möbius-group matrices in the projective space, and as the principal minors of the associated Jacobi matrix. The results have important area of applications to the nonlinear quantum-optical models, including the $k$-photon down-conversion, which play a central role in modern quantum technologies.
@@
where $P_{\mu}$ is the permutation operator acting on the product of $m$ factors, where $\mu$ runs over different orderings of $m$ factors, with $l$ factors of one type and mlm-l of the other type (there are $\binom{m}{l}$ of different orderings). Using Eq. (13) and (14) we get from Eq. (81):